Grade 11IIT JEE Entrance Exam

A car travelling at 60 km/h overtakes another car travelling at 42 km/h. Assuming each car to be 5.0 m long find the time taken during the overtake and the total road distance used for the overtake.

Abhishek Dogra
5 Years agoGrade 11

1 Answer

Rohan Anand
ApprovedApproved Tutor Answer5 Years ago
Relative speed = (60-42) ×5/18
  = 5m/s
Distance = 10m (lenght of both cars)
Time = 10/5 = 2sec
Total distance used = 2×60×(5/18) +5
 = 33.3 + 5
 = 38.3 m 
HOPE IT HELPS!